# Fishers Harvest Parallel Unlearning in Inherited Model Networks

### v1

@@ 3.2. Unified Model Inheritance Graph (UMIG)
Figure 2. The darker blue nodes are all model nodes inheriting from the starting nodes $n_{s}$ and $n_{f}$ in lighter blue. A model node $n_{\ast}$ corresponds to a model $w_{\ast}$.
To uniformly handle and analyze the inheritance and update relationships of models across various learning frameworks, including FL, DDPL, IL, and TL, we propose a DAG-based UMIG. It provides a comprehensive framework to represent components such as models, data, and tasks and their interdependencies within these frameworks. It also facilitates a structured and intuitive understanding of these complex model networks. Figure 2 illustrates the ability of UMIG to depict the inheritance relationships between models.
As illustrated in Figure 2, we define a weighted directional graph $\mathcal{G}=(N,E)$. Herein, $N=\left\{n_{1},n_{2},\ldots,n_{m}\right\}$ represents the model nodes in the network, more specifically, a set of model parameters in a specific state or task. Moreover, $E=\left\{e_{01},e_{02},\ldots,e_{mm}\right\}$ represents the inheritance relationships between model nodes, encompassing parameter transfer and task inheritance. For instance, in FL and DDPL, the edges may represent the transfer of parameters from local to global models. By contrast, in TL, the edges signify the inheritance of tasks from a pre-trained model to a fine-tuned model.
@@
  • Discovery Nodes refer to models trained on data with an unlearned label set, enabling these nodes to adapt to the updated classification ability within the corresponding subgraph of the DAG. The root model nodes $n_{s}$ and $n_{f}$ in Figure 2 belong to this type.
  • Inherited Nodes comprise the remaining nodes in the subgraph, except the discovery nodes. Inherited nodes point to the discovery or other inherited nodes, and focus on refining classification performance. The model nodes $n_{h},n_{j},n_{l},n_{p},n_{g},n_{i},n_{n}$ belong to this type.
3.3. UMIG-Assisted Unlearning
With an emphasis on knowledge inheritance and unlearning tasks, UMIG allows a model to retain and utilize previously learned useful information when receiving new data, achieving an effective combination of old and new knowledge. UMIG also allows the model to remove specific data for continuous optimization and performance improvement.
The UMIG-assisted unlearning consists primarily of the following two processes:
  • Locating unlearning subgraphs. Based on the data that needs to be unlearned, a breadth-first search is conducted to identify the discovery nodes, which then spawn the subgraphs associated with the discovery nodes.
@@ 3.3. UMIG-Assisted Unlearning
By performing a breadth-first search of the entire model network, one or more discovery nodes are identified based on unlearning requests related to specific data. These discovery nodes serve as the roots for the unlearning subgraphs. As depicted in Figure 2, the unlearning subgraphs may overlap since inherited nodes may inherit from multiple discovery nodes. For example, the discovery nodes, $n_{s}$ and $n_{f}$, lead to two overlapping subgraphs with inherited nodes. Overlapping inherited nodes, such as $n_{h},n_{j},n_{l},n_{p}$, would require the elimination of knowledge from multiple discovery nodes when unlearning due to their connections to different unlearning subgraphs. This situation is known as a multi-root scenario. In contrast, inherited nodes within one unlearning subgraph, such as $n_{g},n_{i},n_{n}$ in a single-root scenario, only need to eliminate the impact from a single discovery node.
@@
A complex structure emerges when multiple paths are at varying depths from an inherited node to its corresponding discovery node(s), e.g., $n_{f},n_{g},n_{i}$ or $n_{s},n_{f},n_{h},n_{j}$. If unlearning operations are conducted simultaneously in terms of depth (in other words, nodes at the same depth from the root(s) are processed simultaneously and rely on sequential dependencies), nodes such as $n_{i}$ and $n_{j}$ may require multiple updates. This situation is termed an imbalanced-path scenario, contrasting with a balanced-path scenario (e.g., $n_{f},n_{h},n_{i},n_{l}$), where each node requires only a single update due to uniform path lengths and synchronous processing at each depth. In cases where unlearning is conducted asynchronously, regardless of depth, both scenarios may require multiple updates.
Answer to RQ1: We propose the UMIG to effectively visualize model interdependence and capture the model inheritance in a structured way, leveraging the DAG structure. Assisted by UMIG, the models inheriting the knowledge to be removed can be rapidly identified. Moreover, the UMIG offers a systematic way to collect and prepare inheritance information, thus facilitating the unlearning process of inherited models. We also examine various learning frameworks of inherited models based on the numbers of their upstream models and the paths of knowledge inheritance to elucidate the emerging challenges and considerations of the unlearning process in inherited nodes due to model inheritance.
@@
Figure 3. The unlearning process in $\mathcal{G}$ consists of five steps: 1) Preparation. The discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ are identified by performing a breadth-first search, with unlearning data $\hat{D}^{\scriptscriptstyle D}_{i}$ containing one or more labels. Subsequently, the unlearning graph $\mathcal{\hat{G}}$ is located which root at $\{n^{\scriptscriptstyle D}_{i}\}$. 2) Calculate unlearning the unlearning FIMs $\{\hat{F}_{i}^{\scriptscriptstyle D}\}$ of $\{n^{\scriptscriptstyle D}_{i}\}$. 3) Calculate the model FIM $F_{j}$ of node $n_{j}$ in $\mathcal{\hat{G}}$. 4) Merge unlearning FIMs $\hat{F}^{\scriptscriptstyle D}_{j,k}$ for $n_{j}$ in the case of multiple discovery nodes according to the topology. 5) Update $n_{j}$ by comparing the merged unlearning FIM $\hat{F}_{j}^{\scriptscriptstyle M}$ and $F_{j}$. Steps 2, 3, 4 and 5 apply to all nodes in $\mathcal{\hat{G}}$
4. FIUn-based Parallel Unlearning
@@
This section introduces a novel parallel unlearning method, Fisher Inheritance Unlearning (FIUn), designed to address the challenges associated with machine unlearning in large-scale UMIG model networks. A trusted learning network is considered where all learning contributors follow the designed unlearning process. The FIUn method operates under the general expression provided by the UMIG and leverages FIMs. It initiates the unlearning process by identifying specific model parameters that need updates, making the process more systematic. Derived from models in the unlearning graph, the training data, and the requested unlearning data, the FIMs are crucial in pinpointing essential model parameters and recommending their updated values.
@@
Notably, the FIUn method allows for the independent application of unlearning steps to each impacted model, facilitating a swift and effective parallel unlearning across extensive model networks. The FIUn method is depicted in Figure 3, and the procedural details are given in Algorithm 1.
@@
4Use breadth-first search to identify discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ and consequential unlearning graph $\mathcal{\hat{G}}$
@@
6       Calculate unlearning FIM $\hat{F}^{\scriptscriptstyle D}_{i}$ with Eq. (3)
@@
10       Calculate model FIM $F_{j}$ with Eq. (4)
11       Merge unlearning FIMs $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ with Eq. (5)
12       Update model parameters of node $n_{j}$ with Eq. (6)
@@
4.1. Preparation
Upon receiving an unlearning task $\mathcal{T}$, accompanied by the unlearning dataset $\hat{D}$, FIUn initiates by conducting a breadth-first search from the starting node on the model network $\mathcal{G}$ to identify the discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$. The discovery nodes are the first nodes exhibiting significant recognition of the unlearning data, with $n^{\scriptscriptstyle D}_{i}$ representing the $i$-th discovery node. Subsequently, FIUn delineates the unlearning graph $\mathcal{\hat{G}}$, which is rooted at the identified discovery nodes. FIUn then updates all nodes in the unlearning graph to unlearn the requested unlearning knowledge.
4.2. FIUn
@@
4.2.1. Calculate unlearning FIMs of discovery nodes
The FIUn method initiates by assessing the impact of the unlearning data on the model parameters of the discovery nodes, utilizing the FIM as specified in Eq. (3). This assessment is conducted using the model parameters and the unlearning data (see line 5, Algorithm 1).
@@ 4.2.1. Calculate unlearning FIMs of discovery nodes
(3) $\hat{F}^{\scriptscriptstyle D}_{i}\!\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\right)\!\!\!\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\!\right)^{\!\!\!\intercal}\right)\!\!\Bigg{|}_{w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}}\!\right],$
@@
The unlearning FIM highlights the importance of each parameter in the last layer with respect to the specific unlearning task (Lee et al., 2023; Zhang et al., 2021). In FIUn, the FIM calculation is restricted to the last layer parameters while other layers remain frozen, aimed at reducing the computational complexity associated with FIM calculations, which is a challenge in traditional FIM-based unlearning methods. Using the last layer for unlearning can focus on the most relevant features. Since the last layer directly affects the final decision of the model, adjusting its parameters can effectively facilitate unlearning goals while maintaining stable performance on data not subject to unlearning (Kim et al., 2024).
4.2.2. Calculate model FIMs of all nodes in unlearning graph
The FIUn method proceeds by calculating the model FIMs of all models within the unlearning graph to identify the impact of model parameters in each node (see line 8, Algorithm 1).
For the $j$-th node $n_{j}$ in the unlearning graph $\mathcal{\hat{G}}$, either a discovery node or an inherited node, the model FIM, denoted by $F_{j}$, is given by
(4) $F_{j}\!=\!\mathbb{E}\!\left[\!\left(\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)\!\!\!\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)^{\!\!\!\intercal}\right)\!\Bigg{|}_{w^{\ast}_{D_{j}}}\right]\!,$
@@ 4.2.2. Calculate model FIMs of all nodes in unlearning graph
4.2.3. Merge unlearning FIMs for all nodes in the unlearning graph
@@
The FIUn method estimates the unlearning FIM for all nodes within the unlearning graph, including both discovery and inherited nodes (see line 9, Algorithm 1). The unlearning FIMs from discovery nodes are merged to form a collective unlearning FIM for each node in the unlearning graph. This merger is crucial as the linkages between the unlearning data and parameters are embedded within the model and can persist through model inheritance (Foster et al., 2024b), especially in multi-root scenarios where an inherited node may trace back to several discovery nodes. A novel merging FIM function is introduced, denoted by $\Phi\left(\cdot\right)$, and is defined as
(5) $\hat{F}^{\scriptscriptstyle M}_{j}=\Phi\left(\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}\right),$
@@
where $\hat{F}^{\scriptscriptstyle M}_{j}$ is the merged unlearning FIM for the $j$-th node in the unlearning graph, the set $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ collects the unlearning FIMs of discovery nodes within $\mathcal{\hat{G}}$ that can be traced back from node $n_{j}$, and $k$ denotes the $k$-th discovery node reachable from $n_{j}$. In this paper, the element-wise maximum is employed for the merging FIM function. Specifically, $\hat{F}^{\scriptscriptstyle M}_{j}$ assembles the largest elements among the unlearning FIMs of discovery nodes in multi-root scenarios and takes the only unlearning FIM in single-root scenarios.
@@
Figure 4. There are three discovery nodes that need to undergo the unlearning process, namely model $w_{j}$, model $w_{s}$ and model $w_{f}$. The labels to be unlearned are label 1 and label 2. However, model $w_{j}$ and model $w_{s}$ only have label 1, while model $w_{f}$ only has label 2. Model $w_{h}$ inherits from these three discovery nodes, so we merge the unlearning FIMs of the discovery nodes and perform the unlearning of label 1 and label 2 on model $w_{h}$ together.
@@
As shown in Figure 4, the model $w_{h}$ inherits from models $w_{j}$, $w_{s}$, and $w_{f}$ and needs to remove all requested unlearning data, though each discovery node contains only a portion of the unlearning data. The FIM of each model reflects the sensitivity of its parameters to the unlearning data, which can be interpreted as parameter importance. By taking the maximum of the corresponding elements across these unlearning FIMs, the merged FIM encompasses all unlearning tasks from the discovery nodes and effectively identifies the parameters that need updating to achieve complete unlearning.
4.2.4. Update models in unlearning graph
@@
In the final step, the FIUn method identifies which parameters need to be updated by comparing the merged FIM with the model FIM, as described in line 10 of Algorithm 1. The outcomes of this comparison are also utilized to scale the parameters. For node $n_{j}$ in the unlearning graph, the parameters are updated as
(6) $\hat{w}^{\scriptscriptstyle L}_{j,l}=\begin{cases}\min(\tau\frac{F_{j,l}}{\hat{F}^{\scriptscriptstyle M}_{j,l}},\eta)w^{\scriptscriptstyle L}_{j,l},&\text{if }\frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}>\gamma;\\ w^{\scriptscriptstyle L}_{j,l},&\text{if}\ \frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}\leq\gamma,\end{cases}$
@@
4.3. Discussion

### v4

@@ IV-B Inherited Model Networks
Figure 2: The dark blue nodes are all model nodes inheriting from the starting nodes $n_{s}$ and $n_{f}$ in light blue. A model node $n_{\ast}$ corresponds to a model $w_{\ast}$. $w_{s}$ and $w_{f}$ can originate from different users, with the possibility of simultaneous unlearning requests.
We use a DAG to represent and analyze model inheritance and updates across frameworks like FL, DDPL, IL, and TL. Time is integrated into the representation: each node corresponds to a publication moment, and edges follow the chronological order from left to right. This captures both inheritance dependencies and the temporal sequence of model publication, providing a clear view of complex model networks.
As shown in Fig. 2, we model the system as a weighted directed graph $\mathcal{G}=(N,E)$. The nodes $N={n_{1},n_{2},\ldots,n_{m}}$ represent model states or tasks, each corresponding to a set of parameters. The edges $E={e_{ij}}$ capture inheritance relations, such as parameter transfer or task inheritance. In FL and DDPL, edges may denote parameter transfer from local to global models, while in TL they represent task inheritance from pre-trained to fine-tuned models. Edge direction is determined by temporal order, reflecting when nodes are published, and by reference relations, indicating when one node explicitly refers to another.
@@
  • $\bullet$
    Discovery Nodes are models trained on data with unlearned labels, allowing adaptation to updated classification capabilities within their DAG subgraph. Root model nodes $n_{s}$ and $n_{f}$ exemplify this type in Fig. 2.
  • $\bullet$
    Inherited Nodes are all non-discovery nodes in the subgraph that connect to discovery or other inherited nodes, focusing on classification refinement. Nodes $n_{h},n_{j},n_{l},n_{p},n_{g},n_{i},n_{n}$ belong to this type.
IV-C DAG-Assisted Unlearning
With an emphasis on knowledge inheritance and unlearning tasks, DAG allows a model to retain and utilize previously learned useful information when receiving new data, achieving an effective combination of old and new knowledge. DAG also allows the model to remove specific data for continuous optimization and performance improvement.
The DAG-assisted unlearning consists primarily of the following two processes:
  • $\bullet$
    Locating unlearning subgraphs. When a user wants to unlearn certain data, an unlearn request is initiated. The system first identifies the nodes related to that user, which are referred to as discovery nodes, which then spawn the subgraphs associated with the discovery nodes.
  • $\bullet$
@@ IV-C DAG-Assisted Unlearning
These discovery nodes serve as the roots for the unlearning subgraphs. As depicted in Fig. 2, the unlearning subgraphs may overlap since inherited nodes may inherit from multiple discovery nodes. For example, the discovery nodes, $n_{s}$ and $n_{f}$, lead to two overlapping subgraphs with inherited nodes. Overlapping inherited nodes, such as $n_{h},n_{j},n_{l},n_{p}$, would require the elimination of knowledge from multiple discovery nodes when unlearning due to their connections to different unlearning subgraphs. This situation is known as a multi-root scenario. In contrast, inherited nodes within one unlearning subgraph, such as $n_{g},n_{i},n_{n}$ in a single-root scenario, only need to eliminate the impact from a single discovery node.
@@
A complex structure emerges when multiple paths are at varying depths from an inherited node to its corresponding discovery node(s), e.g., $n_{f},n_{g},n_{i}$ or $n_{s},n_{h},n_{j}$. If unlearning operations are conducted simultaneously in terms of depth (in other words, nodes at the same depth from the root(s) are processed simultaneously and rely on sequential dependencies), nodes such as $n_{i}$ and $n_{j}$ may require multiple updates. This situation is termed an imbalanced-path scenario, contrasting with a balanced-path scenario (e.g., $n_{f},n_{h},n_{i},n_{l}$), where each node requires only a single update due to uniform path lengths and synchronous processing at each depth. In cases where unlearning is conducted asynchronously, regardless of depth, both scenarios may require multiple updates.
Answer to RQ1: We use a DAG to visualize model inheritance. Each node marks a specific time, and edges follow the chronological order of model publication, giving a clear view of complex inheritance networks. The DAG also helps quickly identify models whose knowledge should be removed and systematically collects inheritance information to simplify unlearning. We analyze different learning frameworks, focusing on upstream models and inheritance paths to capture key unlearning scenarios: single-root, multi-root, balanced-path, and imbalanced-path.
@@
Figure 3: The unlearning process in $\mathcal{G}$ consists of five steps: 1) Preparation. The discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ are identified, with unlearning data $\hat{D}^{\scriptscriptstyle D}_{i}$ containing one or more labels. Subsequently, the unlearning graph $\mathcal{\hat{G}}$ is located which root at $\{n^{\scriptscriptstyle D}_{i}\}$. 2) Calculate unlearning the unlearning FIMs $\{\hat{F}_{i}^{\scriptscriptstyle D}\}$ of $\{n^{\scriptscriptstyle D}_{i}\}$. 3) Calculate the model FIM $F_{j}$ of node $n_{j}$ in $\mathcal{\hat{G}}$. 4) Merge unlearning FIMs $\hat{F}^{\scriptscriptstyle D}_{j,k}$ for $n_{j}$ in the case of multiple discovery nodes according to the topology. 5) Update $n_{j}$ by comparing the merged unlearning FIM $\hat{F}_{j}^{\scriptscriptstyle M}$ and $F_{j}$. Steps 2, 3, 4 and 5 apply to all nodes in $\mathcal{\hat{G}}$ .
V FIUn-Based Parallel Unlearning
@@
This section introduces a novel parallel unlearning method, Fisher Inheritance Unlearning (FIUn), designed to address the challenges associated with machine unlearning in large-scale DAG model networks. A trusted learning network is considered where all learning contributors follow the designed unlearning process. The FIUn method operates under the general expression provided by the DAG and leverages FIMs. It initiates the unlearning process by identifying specific model parameters that need updates, making the process more systematic. Derived from models in the unlearning graph, the training data, and the requested unlearning data, the FIMs are crucial in pinpointing essential model parameters and recommending their updated values.
@@
Notably, the FIUn method allows for the independent application of unlearning steps to each impacted model, facilitating a swift and effective parallel unlearning across extensive model networks. The FIUn method is depicted in Fig. 3, and the procedural details are given in Algorithm 1.
@@
4Identify discovery nodes $n^{\scriptscriptstyle D}_{i}$} and the resulting unlearning graph $\mathcal{\hat{G}}$
@@
6  Calculate unlearning FIM $\hat{F}^{\scriptscriptstyle D}_{i}$ with (3)
@@
10  Calculate model FIM $F_{j}$ with (4)
11  Merge unlearning FIMs $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ with (5)
12  Update model parameters of node $n_{j}$ with (6)
@@
V-A FIUn
@@
Preparation. Upon receiving an unlearning task $\mathcal{T}$ initiated by a user, FIUn first identifies the set of discovery nodes $\{n^{\scriptscriptstyle D}_{i}\}$ related to that user from the model network $\mathcal{G}$, with $n^{\scriptscriptstyle D}_{i}$ representing the $i$-th discovery node. Subsequently, FIUn delineates the unlearning graph $\mathcal{\hat{G}}$, which is rooted at the identified discovery nodes. FIUn then updates all nodes in the unlearning graph to unlearn the requested unlearning knowledge.
Step-1: Calculate unlearning FIMs of discovery nodes. The FIUn method initiates by assessing the impact of the unlearning data on the model parameters of the discovery nodes, utilizing the FIM as specified in (3). This assessment is conducted using the model parameters and the unlearning data (see line 5, Algorithm 1).
@@
$\hat{F}^{\scriptscriptstyle D}_{i}\!\!=\!\mathbb{E}\!\left[\!\left(\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\right)\left(\frac{\partial\ln p(\hat{D}_{i}^{\scriptscriptstyle D}|w^{\scriptscriptstyle L}_{i})}{\partial w^{\scriptscriptstyle L}_{i}}\!\right)^{\intercal}\right)\Bigg|_{w^{\ast}_{\hat{D}_{i}^{\scriptscriptstyle D}}}\!\right],$ (3)
@@
The unlearning FIM highlights the importance of last-layer parameters for specific unlearning tasks [lee2023layer]. In FIUn, FIM calculations are restricted to the last layer while freezing other layers, thereby reducing the computational complexity inherent in traditional FIM-based methods. Focusing on the last layer targets the most relevant features, as it directly influences the model’s final decision. This approach effectively facilitates unlearning goals while maintaining stable performance on unaffected data.
Step-2: Calculate model FIMs of all nodes in the unlearning graph. The FIUn method proceeds by calculating the model FIMs of all models within the unlearning graph to identify the impact of model parameters in each node (see line 8, Algorithm 1).
For any node $n_{j}$ in the unlearning graph $\mathcal{\hat{G}}$, the model FIM $F_{j}$ is given by
$F_{j}\!=\!\mathbb{E}\!\left[\!\left(\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)\left(\frac{\partial\ln p(D_{j}|w^{\scriptscriptstyle L}_{j})}{\partial w^{\scriptscriptstyle L}_{j}}\right)^{\intercal}\right)\!\Bigg|_{w^{\ast}_{D_{j}}}\right]\!,$ (4)
@@
Step-3: Merge unlearning FIMs for all nodes in the unlearning graph. The FIUn method estimates the unlearning FIM for all nodes within the unlearning graph, including both discovery and inherited nodes (see line 9, Algorithm 1). The unlearning FIMs from discovery nodes are merged to form a collective unlearning FIM for each node in the unlearning graph. This merger is crucial as the linkages between the unlearning data and parameters are embedded within the model and can persist through model inheritance [foster2024fast], especially in multi-root scenarios of DAG, where an inherited node may trace back to several discovery nodes. Sequential FIM calculation of the unlearned data from multiple roots is required for existing FIM methods [foster2024fast, golatkar2020eternal], but this significantly reduces efficiency in IMNs. To address this, a novel Merging-FIM (MFIM) function is introduced specifically for this context, denoted by $\Phi\left(\cdot\right)$, and is defined as
$\hat{F}^{\scriptscriptstyle M}_{j}=\Phi\left(\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}\right),$ (5)
@@
where $\hat{F}^{\scriptscriptstyle M}_{j}$ is the merged unlearning FIM for the $j$-th node in the unlearning graph, the set $\{\hat{F}^{\scriptscriptstyle D}_{j,k}\}$ collects the unlearning FIMs of discovery nodes within $\mathcal{\hat{G}}$ that can be traced back from node $n_{j}$, and $k$ denotes the $k$-th discovery node reachable from $n_{j}$. In this paper, the element-wise maximum is employed for the MFIM function. Specifically, $\hat{F}^{\scriptscriptstyle M}_{j}$ assembles the largest elements among the unlearning FIMs of discovery nodes in multi-root scenarios and takes the only unlearning FIM in single-root scenarios.
@@
Figure 4: Three discovery nodes need to undergo the unlearning process, namely model $w_{j}$, model $w_{s}$ and model $w_{f}$. The labels to be unlearned are label 1 and label 2. However, model $w_{j}$ and model $w_{s}$ only have label 1, while model $w_{f}$ only has label 2. Model $w_{h}$ inherits from these three discovery nodes, so we merge the unlearning FIMs of the discovery nodes and perform the unlearning of label 1 and label 2 on model $w_{h}$ together.
@@
As shown in Fig. 4, the model $w_{h}$ inherits from models $w_{j}$, $w_{s}$, and $w_{f}$ and needs to remove requested unlearning data, though each discovery node contains only a portion of the unlearning data. The FIM of each model reflects the sensitivity of its parameters to the unlearning data, which can be interpreted as parameter importance. By taking the maximum of the corresponding elements across these unlearning FIMs, the merged FIM encompasses all unlearning tasks from the discovery nodes and identifies the parameters that need updating to achieve complete unlearning in a single process.
@@
Step-4: Update models in the unlearning graph. In the final step, the FIUn method identifies which parameters need to be updated by comparing the merged FIM with the model FIM, as described in line 10 of Algorithm 1. The comparison, based on the Selective Synaptic Dampening (SSD) [foster2024fast], is also utilized to scale the parameters. For node $n_{j}$ in the unlearning graph, the parameters are updated as
$\hat{w}^{\scriptscriptstyle L}_{j,l}=\begin{cases}\min(\tau\frac{F_{j,l}}{\hat{F}^{\scriptscriptstyle M}_{j,l}},\eta)w^{\scriptscriptstyle L}_{j,l},&\text{if }\frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}>\gamma;\\ w^{\scriptscriptstyle L}_{j,l},&\text{if}\ \frac{\hat{F}^{\scriptscriptstyle M}_{j,l}}{F_{j,l}}\leq\gamma,\end{cases}$ (6)
@@
Specific datasets can also be retained by extending the formulations in (5) and (6). In this setting, the unlearning FIM is replaced with a retention FIM, and the role of the parameter $\gamma$ shifts from suppressing to amplifying the weights of critical parameters. Since larger FIM values indicate that the corresponding parameters carry more discriminative information about the training data, we adjust these key parameters proportionally, thereby strengthening the model’s memory and dependence on the retained dataset.
V-B Discussion